# Perseus: Removing Energy Bloat from Large Model Training

### v1

@@ 7                   Remove $i$ from $\mathcal{G}$
       $S$, $T\leftarrow$ FindMinCut($\mathcal{G}$, $p$) $\triangleright$ Modify computation durations ( 5)
       $I\leftarrow\{i\mid i\textrm{ in }S-T\textrm{ cut}\}$ Modify duration of $\forall i\in I$ by $\tau$triangleright Assign frequencies from planned computation times
11       $p^{\prime}\leftarrow$ $\min f_{i}$ that runs no slower than planned return $p^{\prime}$
12
13
14
Algorithm 2 ReduceTime: Reducing the execution time of a computation DAG by $\tau$ with minimal energy increase.
4.4 Extensions
In this section, we present extensions to our optimization algorithm useful for planning large model training.
Tensor Parallelism.
It is trivial to extend our algorithm to tensor parallelism, another essential ingredient of large model training. The observation is that tensor parallel techniques split operations in equal sizes. Therefore, GPUs that execute different portions of the same operation consume the same amount of time and energy. This allows Perseus to operate only on one tensor parallelism copy for each stage, decide the energy schedule for that copy, and replicate it to other copies in the same stage. We show that Perseus works well for hybrid parallelism in Section 6.4.
Single-Choice Operations.
Apart from computation and blocking on communication, there are other operations in the training pipeline that may take non-trivial latencies. For instance, especially on fast GPUs like A100, loading and copying input data into VRAM or communication over slower links can take considerable latency. However, the time and energy consumption of these operations are not affected by the GPU’s frequency. Perseus can take the latency of such operations into account during planning by abstracting them as a node in the DAG with only one frequency choice.
@@ 4.4 Extensions
There are various schedules for pipeline parallel training, including GPipe [26], 1F1B [42], interleaved 1F1B [56], and early recomputation 1F1B [34]. As long as the computations on the schedule can be expressed as a DAG, Perseus can optimize its energy consumption without modification. If there is stage imbalance, we believe that any pipeline schedule will have intrinsic energy bloat.
5 Implementation
The Perseus server is implemented in 1,500 lines of Python [3]. The client is implemented in 300 lines of Python as a library that can be imported into training frameworks.
As a reference, we have integrated the Perseus client with Merak [34], whose open source implementation takes the best of Megatron-LM [5] for high-performance tensor parallelism for transformer models and DeepSpeed [2] for its generic pipeline execution engine. Activation recomputation [12] is enabled by default to allow large batch sizes to fit in GPUs.
@@ 5 Implementation
While accurate timing measurement is well-supported, modern GPUs (NVIDIA A40, A100, and H100) update their cumulative energy consumption counter [8] only once every 100 ms, which is similar to the timescale of one microbatch computation. This makes energy measurement noisy, and no ground truth is available.
At the very moment the cumulative energy counter increases, its value should be accurate. Therefore, Perseus polls the energy counter in the background and records the timestamp of when the counter increased and its value. Then, collected (time, energy) points are interpolated with a linear line. The main training process only records the start and end timing of each computation, allowing Perseus to match timestamps with the interpolated energy curve to calculate each computation’s energy consumption.
The GPU’s frequency is scanned from the highest to the lowest at iteration granularity. As an optimization, when computation energy increases for five consecutive frequencies, profiling is terminated. Beyond that point, frequencies will consume more time and energy, making them suboptimal.
@@
The controller lives in a separate process that communicates with the main training process through a pipe, because setting the GPU’s frequency through NVML [8] should not block computation on the main process. While pipeline execution engine implementations differ widely, many have separate code blocks for forward and backward. Therefore, the controller exposes one method, set_frequency, which takes either "forward" or "backward" and sets the GPU’s frequency as planned.
6 Evaluation
We evaluate Perseus on five workloads and compare it against EnvPipe and Zeus. Our key findings are the following.
@@ 6 Evaluation
  • In emulated large-scale training scenarios, Perseus significantly outperforms the baselines by consistently providing up to 30% energy savings. We also observe that there is a tradeoff between energy savings and scale (§6.3).
@@ 6.1 Experimental Setup
We run our evaluation workloads in a GPU cluster, where each node is equipped with an AMD EPYC 7513 CPU, 512 GB DRAM, and four NVIDIA A40-48G GPUs. For A100 results, we use the node provided by Chameleon Cloud [30], equipped with two Intel Xeon Platinum 8380 CPUs, 512 GB DRAM, and four NVIDIA A100-80G PCIe GPUs.
@@
We evaluate Perseus with various workloads spanning from GPT-3 [11], Bloom [66], BERT [15], T5 [51], to Wide-ResNet [69]. We use model variants with 1.3B to 6.7B parameters to run the models in our testbed, and scale them up to 176B parameters in large-scale emulation. We chose the microbatch size and number of microbatches that yield the highest throughput given the global batch size. We use the minimum imbalance stage partitioning method described in Section 2.2 for all workloads. Appendix A lists complete model configurations, parameters, and stage partitioning details.
@@
We mainly compare against two prior works:
  • EnvPipe [13] reduces only intrinsic energy bloat while trying to minimize slowdown in pipeline execution. We compare Perseus’s amount of energy bloat reduction against that of EnvPipe (§6.2, §6.3).
  • Zeus [68] characterizes the time–energy tradeoff of single GPU training. We compare Perseus’s “iteration time–energy” Pareto frontier against that of Zeus (§6.4).
6.2 Reducing Energy Bloat
We start with overall energy bloat reduction – both intrinsic (§6.2.1) and extrinsic (§6.2.2) – achieved by Perseus and EnvPipe by running various models on our testbed. We compare Perseus against EnvPipe, which reduces energy consumption while trying to minimize slowdown. Both solutions use the same amount of GPU resources.
6.2.1 Intrinsic Energy Bloat Reduction
@@
[…]
Table 2: Intrinsic energy reduction and slowdown comparison between Perseus and EnvPipe.
@@
We make two observations regarding Perseus. First, models show varying amounts of energy savings because their stage imbalance vary (Table 1). For instance, unlike other models, Wide-ResNet 1.5B on A100 after minimum imbalance stage partitioning has nearly perfect stage balance, leaving little room for non-critical computations to slow down.
Second, A40 demonstrates more energy savings compared to A100. This is because the dynamic clock frequency range of A100 (210–1410 MHz) is smaller than that of A40 (210–1740 MHz). Thus, tuning down the GPU’s frequency yields a relatively smaller change in computation time and energy compared to those at the maximum frequency. However, we expect the upcoming NVIDIA H100 GPU to have better savings because its maximum frequency is 1755 MHz for the PCIe version and 1980 MHz for SXM [44].
@@
6.2.2 Intrinsic + Extrinsic Energy Bloat Reduction
@@
[…]
Table 3: Energy savings of Perseus given varying degrees of straggler slowdown ($T^{\prime}/T$). EnvPipe cannot provide additional savings because it does not consider and cannot be applied to optimize extrinsic bloat.
When stragglers create extrinsic energy bloat, the amount of energy savings for a non-straggler pipeline depends on how much energy reduction its time–energy frontier yields for longer iteration times. Table 3 shows the amount of energy savings given varying degrees of straggler slowdown. As described in Section 3.1, non-stragglers’ iteration time is set to be $\min(T^{*},T^{\prime})$, and energy is reduced from (1) slowing down the pipeline itself and (2) blocking on communication for a shorter amount of time, waiting for the straggler. Slowing down the non-stragglers beyond $T^{*}$ increases total energy consumption; hence, Perseus does not slow down pipelines further than $T^{*}$.
The percentage of energy savings increases for lower $T^{\prime}$ values and slowly wanes when $T^{\prime}>T^{*}$. This reduction is due to longer blocking time. That is, the absolute amount of energy savings in Joules is the largest at $T^{*}$, and constant afterward. However, energy consumption during blocking time ($T^{\prime}-T^{*}$) increases for larger $T^{\prime}$ values, lowering the percentage of energy savings for $T^{\prime}>T^{*}$.
Finally, the point of maximum energy savings is different for each model. This is because different models have different $T^{*}$ values, which is determined by how much each stage’s computation slows down on the minimum-energy frequency.
6.3 Large-Scale Emulation
Because we do not have access to a GPU cluster required to run huge models like GPT-3 175B, we use emulation grounded on fine-grained profiling for large-scale evaluation.
Emulation Methodology.
Total
@@ 6.3 Large-Scale Emulation
Table 4: Experiment parameters for GPT-3 175B and Bloom 176B. Each pipeline has tensor parallel degree 8 and 8 pipeline stages.
We profile the time and energy consumption of each layer (e.g., Transformer decoder) in GPT-3 175B and Bloom 176B in bfloat16 to construct the time and energy profile of each stage. Then, we run our optimization algorithm to obtain a theoretical “iteration time–energy” frontier, and use it to report emulated savings. We evaluate the changes in the amount of energy savings in a strong scaling setup presented in Table 4. We do not consider weak scaling, where per-pipeline batch is constant and increasing the global batch size proportionally, because varying the global batch size can affect the model quality [31, 21]. We find that, compared to actual measurements of smaller scale models on A100, the emulator always underestimates the actual percentage of energy savings by 20.2% on average (see Appendix F for details). Meaning, the savings reported by our emulation can likely be considered an empirical lower bound for actual savings. We used A100 SXM GPUs for the A100 results in this section, which is more representative of large scale scenarios (Appendix G provides PCIe results).
@@
Figure 7: Energy savings breakdown of large models emulation with straggler slowdown ($T^{\prime}/T$) 1.20 and 1,024 GPUs.
Figure 7 shows the amount of energy bloat reduction for GPT-3 175B and Bloom 176B large models when slowdown degree is 1.2 on emulated 1,024 GPUs as a representative. EnvPipe can only reduce intrinsic bloat as it does not provide an “iteration time–energy” frontier; even for intrinsic bloat, it reduces less than Perseus. In contrast, Perseus reduces per-iteration energy consumption by up to 30% by reducing both intrinsic and extrinsic energy bloat.
Intrinsic Bloat Reduction.
@@
Table 5: Perseus’s intrinsic energy bloat reduction of each non-straggler pipeline for GPT-3 175B and Bloom 176B with various number of microbatches. All scenarios run with 8 pipeline stages.
@@
Table 5 presents the changes of intrinsic energy bloat saving of a single pipeline in the case of GPT-3 175B and Bloom 176B with respect to various number of microbatches. For all models, as more and more microbatches are added to the pipeline, the amount of intrinsic bloat decreases. This is fundamentally due to the ratio of microbatches in the pipeline’s warm-up and flush phase (beginning and end) vs. steady state phase (middle). Microbatches in the former phase are able to slow down until their minimum energy frequency, yielding large energy savings. However, microbatches in the latter (middle of the pipeline) cannot slow down to their full potential when the amount of stage imbalance is not large, thereby yielding modest savings. When the number of microbatches in the pipeline increases, only the number of steady state microbatches increases, and energy reduction becomes more and more dominated by the average energy savings of steady state microbatches.
Extrinsic Bloat Reduction.
Figure 8: Perseus’s extrinsic energy savings during a single iteration by slowing down non-straggler pipelines. When $T^{\prime}>T^{*}$, slowing non-stragglers down starts increasing energy consumption, hence non-stragglers do not slow down further than $T^{*}$, whose location is denoted by a star. More blocking time after $T^{*}$ decreases overall energy savings. Please note the different Y-axes.
We simulate stragglers in training large models with hybrid parallelism in various strong scaling configurations (Table 4), where a straggler is slower than others in varying degrees from 1.05 to 1.50. Other non-stragglers, after finishing their computation, must wait until the straggler finishes computation, where extrinsic energy bloat emerges. We measure the amount of energy bloat and Perseus’ extrinsic energy savings.
@@
Figure 8 shows the changes of extrinsic energy bloat reduction given varying degrees of straggler slowdown and various strong scale configurations. The trend where energy saving increases until $T^{\prime}<T^{*}$ and wanes afterward, is consistent with what was observed in Section 6.2.2.
@@
An interesting observation here is that there is a tradeoff between scale and energy savings: more pipelines have less percentage of energy savings or less amount of energy savings per pipeline. It may seem intuitive to assume that more pipelines brings more energy savings, as there is only one straggler pipeline that cannot be optimized while all the other pipelines optimize their energy consumption. However, this holds only in weak-scaling configuration, i.e., per-pipeline batch size is constant (increasing the global batch size proportionally to the number of pipelines) and it is not the case in strong-scaling configuration, where the global batch size is constant and per-pipeline batch size is decreased as more pipelines deployed. With less number of microbatches, the ratio of pipeline bubble (time that GPUs are idle) at the beginning and end of each pipeline iteration, which cannot be eliminated by intrinsic energy bloat reduction, increases, resulting in less energy savings.
6.4 Iteration Time–Energy Frontier Comparison

### v2

@@ 12             Remove $i$ from $\mathcal{G}$
18 $S$, $T\leftarrow$ FindMinCut($\mathcal{G}$, $p$)
19
$\triangleright$ Modify computation durations ( 5)
20 Modify duration of $\forall i$ in $S-T$ cut by $\tau$
21
$\triangleright$ Assign frequencies from planned computation times
22 $p^{\prime}\leftarrow$ $\min f_{i}$ that runs no slower than planned
23
24return $p^{\prime}$
25
26
Algorithm 2 GetNextPareto: Reducing the execution time of the DAG by $\tau$ with minimal energy increase.
4.4. Generalizations
In this section, we present generalizations to our optimization algorithm useful for planning large model training.
3D/Hybrid Parallelism.
Operator parallelism techniques (e.g., data, tensor, or sequence parallelism) split operations in equal sizes, resulting in each GPU running the same computation. This allows Perseus to profile only one GPU per stage, decide the energy schedule for that GPU, and replicate it to all other GPUs in the same stage. We show that Perseus works well for 3D parallelism in Section 6.4.
Constant-Time Operations.
Apart from computation and blocking on communication, there are other operations in the training pipeline that may take non-trivial latency. For instance, loading and copying input data into VRAM or communication over slower links can take considerable latency. However, the time and energy consumption of these operations are not affected by the GPU’s frequency. Perseus can take constant-time operations into account during planning by viewing them as a node with only one frequency choice.
@@ 4.4. Generalizations
There are various schedules for pipeline parallel training, including GPipe (gpipe-neurips19, ), 1F1B (pipedreamflush-icml21, ), interleaved 1F1B (megatronlm-arxiv, ), and early recomputation 1F1B (merak-tpds23, ). As long as the computations on the schedule can be expressed as a DAG, Perseus can optimize its energy consumption without modification. As long as there is stage imbalance, any pipeline schedule will have intrinsic energy bloat.
5. Implementation
The Perseus server and client are implemented in Python. Perseus can optimize the energy consumption for any training infrastructure, framework, and accelerator as long as the APIs in Table 2 can be implemented.
As a reference, we have integrated the Perseus client with Merak (merak-tpds23, ), which marries high-performance tensor parallelism of Megatron-LM (megatronlm-github, ) and the generic pipeline execution engine of DeepSpeed (deepspeed-github, ). We provide an example of what is looks like to integrate the client with a training engine in Appendix F. Activation recomputation (checkpointing-arxiv16, ) is enabled to allow large batch sizes to fit in GPUs.
@@ 5. Implementation
The client profiles the time and energy of each forward and backward computation in vivo at the beginning of the training job to take measurements representative of real training. While training engine implementations differ widely, many have separate code blocks for forward and backward, allowing them to be wrapped with the profiler APIs for measurement. The GPU’s frequency is profiled from the highest to the lowest at iteration granularity and stopped when energy consumption increases. Beyond that point, frequencies will consume more time and energy, making them suboptimal.
@@
The client-side controller spawns a separate process that asynchronously sets the GPU’s frequency through NVML (nvml, ) without blocking the main training process. Training frameworks can call set_speed at the beginning of forward or backward to set the GPU’s frequency as planned by the server.
6. Evaluation
We evaluate Perseus on five workloads and compare it against EnvPipe and Zeus. Our key findings are the following:
@@ 6. Evaluation
  • In emulated large-scale training scenarios, Perseus significantly outperforms the baselines by consistently providing up to 30% energy savings (§6.3).
@@ 6.1. Experimental Setup
We run our evaluation workloads in a GPU cluster, where each node is equipped with an AMD EPYC 7513 CPU, 512 GB DRAM, and four NVIDIA A40-48G GPUs. For A100 results, we use a node provided by Chameleon Cloud (chameleoncloud, ), equipped with two Intel Xeon Platinum 8380 CPUs, 512 GB DRAM, and four NVIDIA A100-80G PCIe GPUs.
@@
We evaluate Perseus with various workloads spanning from GPT-3 (gpt3, ), Bloom (bloom-arxiv22, ), BERT (bert, ), T5 (t5-jmlr20, ), to Wide-ResNet (wideresnet-bmvc16, ). We use model variants with 1.3B to 6.7B parameters to run the models in our testbed, and scale them up to 176B parameters in large-scale emulation. We chose the microbatch size and number of microbatches that yield the highest throughput given the global batch size. We use the minimum imbalance stage partitioning method described in Section 2.2 for all workloads. Appendix B lists complete model configurations, parameters, and stage partitioning details.
Metrics.
We report energy reduction and slowdown of a training iteration (%), relative to using all maximum GPU frequencies. In most cases slowdown is close to zero, in which case energy and power reductions coincide. Reducing only extrinsic bloat is not possible, because Perseus reduces extrinsic bloat while keeping intrinsic bloat low as it slows down non-straggler pipelines. Therefore, we report (1) intrinsic bloat reduction without stragglers and (2) intrinsic + extrinsic bloat reduction with stragglers.
@@
We mainly compare with two prior works:
  • EnvPipe (envpipe-atc23, ) reduces only intrinsic energy bloat while trying to minimize slowdown. We compare Perseus’s energy bloat reduction with EnvPipe (§6.2, §6.3).
  • Zeus (zeus-nsdi23, ) characterizes the time–energy tradeoff of single GPU training. We compare Perseus’s “iteration time–energy” Pareto frontier against that of Zeus (§6.4).
6.2. Reducing Energy Bloat
We start with overall energy bloat reduction – intrinsic bloat without stragglers (§6.2.1) and intrinsic + extrinsic bloat with stragglers (§6.2.2) – achieved by Perseus and EnvPipe. All numbers were obtained by running on testbed GPUs, and all solutions use the same amount of GPU resources.
6.2.1. Intrinsic Bloat Reduction Without Stragglers
@@
(b) Eight stage pipeline parallelism on A40 GPUs
Table 3. [Experiment] Intrinsic energy bloat (without stragglers) reduction and iteration time slowdown.
@@
We make two observations regarding Perseus. First, models show varying amounts of energy savings because (1) their stage imbalance vary (Table 1), and (2) their forward and backward are composed on different computations, which affects time/energy sensitivity when changing the frequency.
Second, A40 demonstrates more energy savings compared to A100. This is because the dynamic clock frequency range of A100 (210–1410 MHz) is smaller than that of A40 (210–1740 MHz). Thus, tuning down the GPU’s frequency yields a relatively smaller change in computation time and energy compared to those at the maximum frequency. However, we expect the more recent GPUs to have better percentage savings due to higher maximum frequency (e.g., 1980 MHz for H100 SXM (H100whitepaper, )) and absolute savings due to high TDP (e.g., 1,200 W for each GPU on GB200 (blackwell-news, )).
@@
6.2.2. Intrinsic + Extrinsic Bloat Reduction With Stragglers
@@
(b) Eight stage pipeline parallelism on A40 GPUs
Table 4. [Experiment] Energy savings given varying straggler slowdown ($T^{\prime}/T$). Perseus can reduce extrinsic bloat while keeping intrinsic bloat low, whereas EnvPipe cannot.
When stragglers create extrinsic energy bloat, the amount of energy savings for a non-straggler pipeline depends on how much energy reduction its “iteration time–energy” Pareto frontier yields for longer iteration times. Table 4 shows the amount of energy savings of a pipeline given varying degrees of straggler slowdown. As described in Section 3.1, a non-straggler pipeline’s iteration time is set to be $T_{\textrm{opt}}=\min(T^{*},T^{\prime})$, and energy reduction comes from (1) slowing down the pipeline itself and (2) reducing the time blocking on communication, waiting for the straggler. Slowing down the non-stragglers beyond $T^{*}$ will increase total energy consumption; hence, Perseus does not slow down pipelines further than $T^{*}$.
The percentage of savings first increases with the straggler’s iteration time, but then slowly wanes as it slows down beyond $T^{*}$. This is expected. The absolute amount of energy reduction in Joules is the largest when the non-straggler pipeline’s iteration time is $T^{*}$ and constant afterward, because Perseus does not slow down non-straggler pipelines beyond $T^{*}$. Thus, as the straggler slows down beyond $T^{*}$, the time and energy consumed while waiting for the straggler increases, lowering the percentage of energy savings.
Finally, the point of maximum energy savings is different for each model. This is because each model has a different $T^{*}$ value, which is determined by how much each stage’s computation slows down on the minimum-energy frequency.
6.2.3. How Much Potential Savings Was Realized?
The largest possible savings under our problem setting occurs when running every computation at their minimum-energy frequencies (i.e., the $T^{*}$ point on the “iteration time–energy” Pareto frontier). For intrinsic bloat without stragglers, Perseus on average realizes 74% and 89% the opportunity on A100 and A40, respectively, with negligible slowdown. This is possible because there are much more non-critical computations in the DAG that can be slowed down than critical ones. With stragglers, Perseus fully realizes potential savings when the straggler’s slowdown degrees are on average 1.1 and 1.15 on A100 and A40 respectively, which is not unrealistic considering slowdowns reported in literature (§2.3).
6.3. Large-Scale Emulation
Because we do not have access to a GPU cluster required to run huge models like GPT-3 175B, we use emulation grounded on fine-grained profiling for large-scale evaluation. In general, trends in our emulation result match those obtained from real training in Section 6.2.
@@ 6.3. Large-Scale Emulation
Table 5. Strong scaling parameters for large-scale emulation. A pipeline has tensor parallel degree 8 and 8 pipeline stages.
Emulation Methodology.
We profile the time and energy consumption of each layer (e.g., Transformer decoder) in GPT-3 175B and Bloom 176B and run our optimization algorithm to obtain the “iteration time–energy” frontier. We perform strong scaling when varying the number of GPUs (Table 5) in order to keep the global batch size constant (largebatch1, ; largebatch2, ). We used A100 SXM GPUs for emulation, which we believe are more representative of large-scale training infrastructure.
Emulator Fidelity.
We compare the percentage of energy savings vs. running all maximum frequencies given by our emulator and real experiments for our A100 workloads and find that the emulator always underestimates energy savings. Specifically, savings on the leftmost and rightmost point of the Pareto frontier are underestimated by 18.6% and 21.7% on average, respectively. We believe this is due to our simplifying assumption that $P_{\mathrm{blocking}}$ is constant regardless of the GPU’s frequency. This means savings given by the emulator can be considered a lower bound for actual savings.
Figure 7. [Emulation] Energy savings breakdown with straggler slowdown 1.2 and 1,024 GPUs.
@@
Figure 7 breaks down the amount of energy bloat reduction for GPT-3 175B and Bloom 176B when slowdown degree is 1.2 on 1,024 GPUs. EnvPipe can only reduce intrinsic bloat as it does not provide an “iteration time–energy” frontier; even for intrinsic bloat, it is suboptimal. In contrast, Perseus reduces energy consumption by up to 30% by reducing both intrinsic and extrinsic energy bloat.
@@
Table 6. [Emulation] Perseus’s intrinsic energy bloat reduction without stragglers for GPT-3 175B and Bloom 176B. Number of microbatches is varied following Table 5.
Intrinsic Bloat Reduction Without Stragglers.
@@
Table 6 shows Perseus’s intrinsic energy bloat reduction without stragglers for GPT-3 175B and Bloom 176B. The number of microbatches was varied based on Table 5. For all models, as more and more microbatches are added to the pipeline, the amount of intrinsic bloat decreases. This is fundamentally due to the ratio of microbatches in the 1F1B’s warm-up and flush phase (beginning and end) vs. steady state phase (middle). Most microbatches in the warm-up and flush phases can slow down until their minimum energy frequency, yielding large energy savings. However, microbatches in the pipeline’s steady state cannot slow down to their full potential when the amount of stage imbalance is not large, thereby yielding modest savings. When the number of microbatches in the pipeline increases, only the number of steady state microbatches increases, and energy reduction converges to the average energy savings of steady state microbatches.
Figure 8. [Emulation] Perseus’s intrinsic + extrinsic energy bloat reduction with varying straggler slowdown ($T^{\prime}/T$). Number of pipelines is varied following Table 5. $\bigstar$ denotes $T^{*}/T$ for each pipeline. Please note the different Y-axes.
Intrinsic + Extrinsic Bloat Reduction With Stragglers.
@@
We introduce stragglers in large-scale emulation in various strong scaling configurations (Table 5) and with varying straggler slowdown degree. Figure 8 reports the amount of intrinsic + extrinsic energy bloat reduction achieved by Perseus. The trend where energy saving increases until $T^{\prime}<T^{*}$ and wanes afterward is consistent with Section 6.2.2.
@@
An interesting observation here is that there is a tradeoff between scale and energy savings: configurations with more pipelines have less percentage of energy savings or less amount of energy savings per pipeline. It may seem intuitive to assume that more pipelines brings more energy savings, as there is only one straggler pipeline and all the other pipelines can reduce their energy consumption. However, this holds only in weak scaling configuration, where per-pipeline batch size is held constant (i.e., increasing the global batch size proportionally to the number of pipelines). Instead, in the more realistic strong scaling configuration (i.e., the global batch size is constant and per-pipeline batch size is decreased as more pipelines deployed), each pipeline’s number of microbatches changes. With fewer microbatches, the ratio of pipeline bubble (time that GPUs are idle) at the beginning and end of each pipeline iteration increases (megatronlm-sc21, ). These bubbles cannot be perfectly eliminated by intrinsic energy bloat reduction, resulting in a smaller energy saving percentages. However, as the absolute number of GPUs increase, even a small savings percentage gives huge absolute energy savings.
6.4. Iteration Time–Energy Frontier Comparison